If a die is rolled 360 times, find the standard deviation for the number of 3s that will be rolled.
Answer by stanbon(75887)

If a die is rolled 360 times, find the standard deviation for the number of 3s that will be rolled.
----------
standard deviation = sqrt(npq)
= sqrt(360*(1/6)(5/6)) = 7.0711
---------
Cheers,
Stan H.